Fred invested $45,000 for one year, part 5% and part at 3%. if he earned a total of $1910 for the year, how much was invested at each rate?
Let x=

Answer by ReadingBoosters(3246)

x = $ at 5%
y = $ at 3%
x+y = 45000
.05x + .03y = 1910
.05(45000-y) + .03y = 1910
2250 - .05y + .03y = 1910
2250 - .02y = 1910
-.02y = -340
y = $17,000 at 3%
x = 45000 - 17000 = $28,000 at 5%
Proof
.05(28000)+.03(17000) = 1400 + 510 = 1910
..................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.